(a) Set up a table of microstates to show that the ground term for the d1 ion is the singlet 2D. What are the components of this term in a tetrahedral field?

(b) Repeat the process for a d2 ion and show that the ground and excited terms are 3F and 3P. What are the components of these terms in tetrahedral and octahedral fields?
Answer #1


Sak 42 => 25+1= 2x4+1= 2 L=0.1234: sple FG Term Symbol + 25+ L3 Ja tj 17 Now I This is ground Term Symbol becous. I have Lowe micro states. a), that @ d » 20 ico State => L - 40 - I 1037 microstates = 10 I l microstate components in Tofield - /2T2 7 ETerm for ground Symbol =) 25+127 - Saltdal =) L= 2+1 423 F 25+1=3 =) J= Lts.... rs => 4 3 2 = 3 F4 3 F3 (3) ground Term >> lo6 = te qe J2 3F and 3p. microstats Llo = L [2110-2 218 mice state =) 45 Id2 miuostatel octahydral | 3F | 45 | 31- K 312g F L